Solve the equation by completing the square
x^2 + 6x − 12 = 0

Answer by richard1234(7193)

Rewrite as (x+3)^2 - 21 = 0.

(x+3)^2 = 21

x+3 = +/- sqrt(21)

x = +/- sqrt(21) - 3